# Quadratic Motion Polynomials With Irregular Factorizations

### v1

@@ Abstract
There exists an algorithm for the factorization of motion polynomials that works in generic cases. It hinges on the invertibility of a certain coefficient occurring in the algorithm. If this coefficient is not invertible, factorizations may or may not exist. In the case of existence we call this an irregular factorization. We characterize quadratic motion polynomials with irregular factorizations in terms of algebraic equations and present examples whose number of unique factorizations range from one to infinitely many. For two special sub-cases we show the unique existence of such polynomials. In case of commuting factors we obtain the conformal Villarceau motion, in case of rigid body motions the circular translation.
@@ 1 Introduction
While generically a motion polynomial of degree $n$ admits $n!$ factorizations with linear factors over both, the quaternions $\mathbb{H}$ and the dual quaternions $\mathbb{D}\mathbb{H}$, a notable difference between these two algebras is that motion polynomials over $\mathbb{D}\mathbb{H}$ might also have infinitely many or no factorization. While the case of zero factorizations has been resolved recently [8], not much is known about motion polynomials with infinite factorizations. The most basic example is a quadratic motion polynomial that parametrizes the curvilinear translation along a circle (a circular translation for short). We will re-visit it in Section 4. The kinematic explanation for its unusual factorization properties is the possibility to generate this motion in infinitely many ways by a parallelogram linkage, as seen in Figure 1.
@@
The concept of motion polynomials [^1] We finally settled with this name after calling them “rotor polynomials” in the proposal acknowledged at the end of this text and “spinor polynomials” in [9]. Both earlier names led to confusion in relevant scientific communities. and the factorization algorithm of [1] readily extends from dual quaternions $\mathbb{D}\mathbb{H}$ to conformal geometric algebra $\mathrm{CGA}$ or from Euclidean kinematics to conformal kinematics [10]. The most notable difference is that a generic motion polynomial of degree $n$ generically admits up to $(2n)!/2^{n}$ factorizations. Even in this larger algebra, only one further non-trivial motion polynomial with infinite factorizations is known (cf. Definition 4.1). Motivated by some applications to physics, it was introduced by L. Dorst in [11]. Its infinitely many factorizations were described in [12] and it is noteworthy that they all commute. Since the motion’s trajectories are related to Villarceau circles on a torus, as can be observed in Figure 2 , we call it the conformal Villarceau motion, cf. Section 5.
@@
The primary approach involves examining the conditions under which an algebra element becomes non-invertible. The characteristics of non-invertible vectors in Conformal Geometric Algebra ($\mathrm{CGA}$) have been extensively investigated in the literature [13, 14]. These findings have significant applications in the field of Automated Theorem Proving, as demonstrated in several studies [15, 16, 17]. This paper extends the investigation to non-invertible elements that go beyond merely null vectors.
@@ 2.1 Conformal Geometric Algebra
To describe conformal motions we will use the framework of conformal geometric algebra also known as $\mathrm{CGA}$ [18]. It is a Clifford-algebra of signature $(4,1)$ over the reals together with an involution $a\mapsto\tilde{a}$ called reversion. We choose the orthonormal basis $\{e_{1},e_{2},e_{3},e_{+},e_{-}\}\in\mathbb{R}^{4,1}$ such that
@@
For $i\neq j\in\{1,2,3,+,-\}$ we define $e_{i}e_{j}=-e_{j}e_{i}\coloneqq e_{ij}.$ This can be extended in the same way to the product of multiple vectors. The elements of $\mathrm{CGA}$ consist of linear combinations of all possible multiplications of the basis vectors. We say an element has grade $n$ if it can be written as the product of $n$ vectors, i.e. $n$ linear combinations of the basis elements. In this case we call it an $n$-vector. If an element does not have a unique grade, but is rather the sum of multiple elements with a single grade, we call it a multivector.
@@
A conformal displacement is the successive inversion in a number of spheres. The sphere with center $(c_{x},c_{y},c_{z})$ and radius $r$ is embedded in $\mathrm{CGA}$ as
$s\coloneqq c_{x}e_{1}+c_{y}e_{2}+c_{z}e_{3}+\frac{q-1}{2}e_{+}+\frac{q+1}{2}e_{-},$
where q := $c_{x}^{2}+c_{y}^{2}+c_{z}^{2}-r^{2}$.
Points and planes are viewed as special cases of spheres. For points we let the radius be zero and planes can be thought of taking the limit of the center and the radius going to inifinity. From this it follows that points at $(p_{x},p_{y},p_{z})$ and planes with normal vector $(n_{x},n_{y},n_{z})$ and distance $d$ to the origin are represented as
$p\coloneqq p_{x}e_{1}+p_{y}e_{2}+p_{z}e_{3}+\frac{p_{x}^{2}+p_{y}^{2}+p_{z}^{2}-1}{2}e_{+}+\frac{p_{x}^{2}+p_{y}^{2}+p_{z}^{2}+1}{2}e_{-},$
@@
$pl\coloneqq n_{x}e_{1}+n_{y}e_{2}+n_{z}e_{3}+(n_{x}^{2}+n_{y}^{2}+n_{z}^{2})d(e_{+}+e_{-}),$
The inversion of an element $a$ by a sphere $s$ is given by $sa\tilde{s}$, commonly known as the sandwich product. Note that we do not require $s\tilde{s}=\pm 1$, mostly for ease of notation in the context of motion polynomials. Since we do not care about normalization, a projective viewpoint is often more natural and we will often consider spheres and other geometric entities as points of the projective space $\mathbb{P}(\mathrm{CGA})$ over $\mathrm{CGA}$. If $a\in\mathrm{CGA}$ then we denote the corresponding point in $\mathbb{P}(\mathrm{CGA})$ by $[a]$.
@@
An element $[a]\in\mathbb{P}(\mathrm{CGA}_{+})$ describes a conformal displacement if and only if $a\tilde{a}$, $\tilde{a}a\in\mathbb{R}\setminus\{0\}$. Elements $[a]\in\mathbb{P}(\mathrm{CGA}_{+})$ fulfilling the condition $a\tilde{a}=\tilde{a}a\in\mathbb{R}$ lie on an algebraic variety defined by it, called the Study variety $\mathcal{S}$ of conformal kinematics [9].
Following a suggestion by N. Wildberger [19], we call $a\tilde{a}$ the quadrance of $a$ if $a\tilde{a}\in\mathbb{R}$. Sometimes this is also called the norm of $a$, but since it is rather a squared norm we chose the former name to avoid confusion. Note that we do not have to distinguish a left and right quadrance $a\tilde{a}$, $\tilde{a}a$ since both values coincide.
To describe smooth motions we can use the typical approach of taking the rotor exponential. In the projective setting the exponential $e^{ua}$ can be rewritten via a real non-linear reparametrization of $u$ as a linear polynomial function $t-a$. For $a\in\mathrm{CGA}_{+}$ we call $t-a$ a simple motion if $(t-a)\widetilde{(t-a)}\in\mathbb{R}$ for any $t\in\mathbb{R}$. These can be categorized into three groups according to the number of distinct real roots of the quadrance-polynomial $(t-a)\widetilde{(t-a)}$. For zero roots it describes a rotation, for one distinct root a translation and for two a scaling [13, 9].
@@ 2.2 Polynomials in $\mathrm{CGA}_{+}$
We will now take a closer look at polynomials in $\mathrm{CGA}_{+}$. Let $C=\sum_{i=0}^{n}c_{i}t^{i}$, where $c_{i}\in\mathrm{CGA}_{+}$ be a polynomial in the indeterminate $t$. We define the multiplication with the convention that the indeterminate commutes with the coefficients of the polynomials. This is reasonable given the fact the we regard $t$ as a real motion parameter. Nonetheless, we will also evaluate polynomials at more general algebra elements, but then we need to distinguish two types of evaluation, the right and left evaluation. They are defined by
$C(h)_{r}\coloneqq\sum_{i=1}^{n}c_{i}h^{i}\quad\text{and}\quad C(h)_{l}\coloneqq\sum_{i=1}^{n}h^{i}c_{i},$
@@
We define the reversion $\tilde{C}$ of $C$ by taking the reversion of all coefficients. Furthermore we define the action of $C$ on an element $a\in\mathrm{CGA}$ by the sandwich product $Ca\tilde{C}$. For $C$ to describe a conformal motion it is necessary for its left and right quadrance to be equal and a non-zero real polynomial: $C\tilde{C}=\tilde{C}C\in\mathbb{R}[t]\setminus\{0\}$. In the case that these conditions are met we call $C$ a motion polynomial. Indeed, it describes a conformal motion as $C(t)$ is a conformal displacement for any $t\in\mathbb{R}$ with at most finitely many exceptions, namely the real roots of the quadrance polynomial $C\tilde{C}$. Since the trajectories of all points are rational curves we speak of a rational motion.
@@ 2.3 Decomposition into Simple Motions
    For generic polynomials $C$ of degree $n$ the amount of factorizations depends upon the number of real roots of $C\tilde{C}$ and ranges from $n!$ for no real roots and $\frac{(2n)!}{2^{n}}$ for $2n$ roots.
@@
    $t-h$ is a right factor of $C$ if and only if $h$ is a right root of $C$
@@
For this we will henceforth assume $C$ to have an invertible leading coefficient. Because of $C\tilde{C}\neq 0$ this is no loss of generality as it can be ensured by a suitable rational re-parametrization. But then we can also assume that $C$ is monic as the leading coefficient can be factored out.
Since $t-h$ being a right factor is equivalent to $h$ being a right root of $C$ the question of finding right factors reduces to the question of finding right roots. Since a right root of the polynomial is also necessarily a right root of the quadrance of the polynomial, we can search for roots of $C\tilde{C}$. Each right factor $t-h$ and therefore also each right root corresponds to a monic, quadratic factor $M\coloneqq(t-h)\widetilde{(t-h)}$ of the quadrance polynomial. Using the Euclidean algorithm [10] (This is possible because the leading coefficient $1$ of the divisor $t-h$ is invertible.) we can now divide $C$ by $M$ and get
@@
for suitable polynomials $Q$, $R$ with $\deg R<\deg M=2$. Since we know that $h$ is a root of $C$ and $M$ it follows that it also has to be a root of $R$ [1, Lemma 1]. Furthermore we can write $R=r_{1}t+r_{0}$. The question of factorization is closely related to roots of the linear polynomial $R$. Assuming for the moment that $r_{1}$ is invertible we get the unique solution $h=-r_{1}^{-1}r_{0}$ for our particular choice of $M$. Once we have found this right factor we can divide $C$ by it and get $C^{\prime}$ of a lower degree and can start from the top. Note that there also a “left” variant of this factorization algorithm.
If the leading coefficient $r_{1}$ is not invertible, we potentially get no or infinite roots of $R$. To solve for $h$ we convert $R(h)=0$ into a linear system of equations via the coefficient-vectors with regards to the basis of $\mathrm{CGA}$. The resulting system of linear equations has no or an infinite amount of solutions which may lead to no, or infinitely many but also to finitely many factors of $C$ that can be determined by imposing further necessary condition (cf. Section 3.3). In order to capture this special situation, we define
@@
We call a factorization $C=(t-h_{1})(t-h_{2})\cdots(t-h_{n})$ of a motion polynomial $C$ irregular, if there exists an index $\ell\in\{1,2,\ldots,n\}$ such the linear remainder polynomial $R$ obtained by dividing either
@@
by $M\coloneqq(t-h_{\ell})(t-\tilde{h}_{\ell})$ has a non-invertible leading coefficient. In this case we also say that $C$ is irregularly factorizable.
@@
Definition 2.1 covers both, the left and the right version of the factorization algorithm. It is, however, not clear whether an irregular “right” factorization implies an irregular “left” factorization or not. For quadratic motion polynomials however, both notions coincide.

### v2

@@ Abstract
Motion polynomials are a specific type of polynomial over a Clifford algebra that can conveniently describe rational motions. There exists an algorithm for the factorization of motion polynomials that works in generic cases. It hinges on the invertibility of a certain coefficient occurring in the algorithm. If this coefficient is not invertible, factorizations may or may not exist. In the case of existence we call this an irregular factorization. We characterize quadratic motion polynomials with irregular factorizations in terms of algebraic equations and present examples whose number of unique factorizations range from one to infinitely many. For two special sub-cases we show the unique existence of such polynomials. In case of commuting factors we obtain the conformal Villarceau motion, in case of rigid body motions the circular translation.
@@ 1 Introduction
While generically a motion polynomial of degree $n$ admits $n!$ factorizations with linear factors over both the quaternions $\mathbb{H}$ and the dual quaternions $\mathbb{DH}$, a notable difference between these two algebras is that motion polynomials over $\mathbb{DH}$ might also have infinitely many or no factorization. While the case of zero factorizations has been resolved recently [8], not much is known about motion polynomials with infinite factorizations. The most basic example is a quadratic motion polynomial that parametrizes the curvilinear translation along a circle (a circular translation for short). We will re-visit it in Section 4. The kinematic explanation for its unusual factorization properties is the possibility to generate this motion in infinitely many ways by a parallelogram linkage, as seen in Figure 1.
@@
The concept of motion polynomials [^1] We finally settled with this name after calling them “rotor polynomials” in the proposal acknowledged at the end of this text and “spinor polynomials” in [9]. Both earlier names led to confusion in relevant scientific communities. and the factorization algorithm of [1] readily extends from dual quaternions $\mathbb{DH}$ to conformal geometric algebra $\mathrm{CGA}$ or from Euclidean kinematics to conformal kinematics [10]. The most notable difference is that a generic motion polynomial of degree $n$ generically admits up to $(2n)!/2^{n}$ factorizations. Even in this larger algebra, only one further non-trivial motion polynomial with infinite factorizations is known (cf. Definition 4.1). Motivated by some applications to physics, it was introduced by L. Dorst in [11]. Its infinitely many factorizations were described in [12] and it is noteworthy that they all commute. Since the motion’s trajectories are related to Villarceau circles on a torus, as can be observed in Figure 2, we call it the conformal Villarceau motion, cf. Section 5.
@@
The primary approach involves examining the conditions under which an algebra element becomes non-invertible. The characteristics of non-invertible vectors in Conformal Geometric Algebra ($\mathrm{CGA}$) have been extensively investigated in the literature [13, 14]. These findings have significant applications in the field of Automated Theorem Proving, as demonstrated in several studies [15, 16, 17]. This paper extends the investigation to non-invertible elements that go beyond mere null vectors.
@@ 2.1 Conformal Geometric Algebra
To describe conformal motions, which contain the Euclidean motions as a subset, we will use the framework of conformal geometric algebra also known as $\mathrm{CGA}$ [18]. It is a Clifford algebra of signature $(4,1)$ over the reals together with an involution $a\mapsto\tilde{a}$ called reversion. We choose the orthonormal basis $\{e_{1},e_{2},e_{3},e_{+},e_{-}\}\in\mathbb{R}^{4,1}$ such that
@@
For $i\neq j\in\{1,2,3,+,-\}$ we define $e_{i}e_{j}=-e_{j}e_{i}\coloneqq e_{ij}.$ This can be extended in the same way to the product of multiple basis vectors. The elements of $\mathrm{CGA}$ consist of linear combinations of all possible multiplications of the basis vectors. We say an element has grade $n$ if it can be written as a linear combination of products of $n$ basis vectors. In this case we call it an $n$-vector. If an element does not have a unique grade, but is rather the sum of multiple elements with a single grade, we call it a multivector.
@@
For our approach, a projective viewpoint is more natural and we will often consider multivectors as points of the projective space $\mathbb{P}(\mathrm{CGA})$ over $\mathrm{CGA}$. If $a\in\mathrm{CGA}$ then we denote the corresponding point in $\mathbb{P}(\mathrm{CGA})$ by $[a]$.
A conformal displacement is a successive inversion in a number of spheres. The sphere $s$ with center $(c_{x},c_{y},c_{z})$ and radius $r$ is embedded in $\mathrm{CGA}$ as
$s=c_{x}e_{1}+c_{y}e_{2}+c_{z}e_{3}+\frac{q-1}{2}e_{+}+\frac{q+1}{2}e_{-},$
where $q\coloneqq c_{x}^{2}+c_{y}^{2}+c_{z}^{2}-r^{2}$.
Points and planes are viewed as special cases of spheres. For points we let the radius be zero and planes can be thought of taking the limit of the center and the radius going to inifinity. From this it follows that the point $p$ at $(p_{x},p_{y},p_{z})$ and the plane $\pi$ with normal vector $(n_{x},n_{y},n_{z})$ and distance $d$ to the origin are represented as
$p=p_{x}e_{1}+p_{y}e_{2}+p_{z}e_{3}+\frac{p_{x}^{2}+p_{y}^{2}+p_{z}^{2}-1}{2}e_{+}+\frac{p_{x}^{2}+p_{y}^{2}+p_{z}^{2}+1}{2}e_{-},$
@@
$\pi=n_{x}e_{1}+n_{y}e_{2}+n_{z}e_{3}+(n_{x}^{2}+n_{y}^{2}+n_{z}^{2})d(e_{+}+e_{-}).$
The inversion of an element $[a]$ by a sphere $s$ is given by $sa\tilde{s}$], commonly known as the sandwich product. As already mentioned, using equivalence classes is beneficial in the context of motion polynomials. In this viewpoint we loose the weight and orientation of objects but gain ease of use for polynomial computations. Instead of the sphere $s$ we could equally well use $[s]$. Hence we also do not require the normalization $s\tilde{s}=\pm 1$.
@@
[…]
Following a suggestion by N. Wildberger [19], we call $a\tilde{a}$ the quadrance of $a$ if $a\tilde{a}\in\mathbb{R}$. Sometimes this is also called the norm of $a$, but since it is rather a squared norm we chose the former name to avoid confusion. Note that we do not have to distinguish a left and right quadrance $\tilde{a}, a~a\tilde${a}a since both values coincide.
To describe smooth motions we can use the typical approach of taking the rotor exponential. Let $B\in\mathrm{CGA}$ be a 2-blade. The motion described by this blade and parametrized by a time $u$ is then given by sandwiching with the exponential $e^{uB/2}$. Using the fact that
$e^{uB^{\prime}}=\left\{\begin{array}[]{l}\cos(u\sqrt{B^{\prime}\tilde{B}^{\prime}})+\frac{B^{\prime}}{\sqrt{B^{\prime}\tilde{B}^{\prime}}}\sin(u\sqrt{B^{\prime}\tilde{B}^{\prime}})\\ 1+B^{\prime}u\\ \cosh(u\sqrt{-B^{\prime}\tilde{B}^{\prime}})+\frac{B^{\prime}}{\sqrt{-B^{\prime}\tilde{B}^{\prime}}}\sinh(u\sqrt{-B^{\prime}\tilde{B}^{\prime}})\end{array}\right.\begin{array}[]{c}\text{ if }B^{\prime}\tilde{B}^{\prime}>0\\ \text{ if }B^{\prime}\tilde{B}^{\prime}=0\\ \text{ if }B^{\prime}\tilde{B}^{\prime}<0\end{array},$
we can rewrite the exponential $e^{uB/2}$, as a linear polynomial function tat-a via an appropriate choice of the following reparametrizations.
$t\coloneqq\cot(u),\quad t\coloneqq u^{-1},\quad t\coloneqq\pm\coth(u)$
For $a\in\mathrm{CGA}_{+}$ we call tat-a a simple motion if $(t-a)(t-a){\,\tilde{}}\in\mathbb{R}$ for any $t\in\mathbb{R}$. These can be categorized into three groups according to the number of distinct real roots of the quadrance-polynomial $(t-a)(t-a){\,\tilde{}}$. For zero roots it describes a rotation, for one distinct root a translation and for two a scaling [13, 9].
The number of roots correspond to the number of distinct points of a transformation to which all points in space are mapped simultaneously, as seen in Figure 3. For a rotation there exists no such point. For a translation this point is the point at infinity, or a conformal image thereof. For a scaling the points are the center of scaling and the point at infinity, or their conformal images.
Figure 3: Simple conformal motions and euclidean counterpart. From left to right: rotation, transversion, scaling
@@ 2.2 Polynomials in $\mathrm{CGA}_{+}$
We will now take a closer look at polynomials in $\mathrm{CGA}_{+}$. Let $C=\sum_{i=0}^{n}c_{i}t^{i}$, where $c_{i}\in\mathrm{CGA}_{+}$ be a polynomial in the indeterminate $t$. We define the multiplication with the convention that the indeterminate commutes with the coefficients of the polynomials. This is reasonable given the fact the we regard $t$ as a real motion parameter (time). Nonetheless, we will also evaluate polynomials at more general algebra elements, but then we need to distinguish two types of evaluation, the right and left evaluation. They are defined by
$C(h)_{r}\coloneqq\sum_{i=0}^{n}c_{i}h^{i}\quad\text{and}\quad C(h)_{l}\coloneqq\sum_{i=0}^{n}h^{i}c_{i},$
@@
We define the reversion $\tilde{C}$ of $C$ by taking the reversion of all coefficients. Furthermore, we define the action of $C$ on an element $a\in\mathrm{CGA}$ by the sandwich product $Ca\tilde{C}$. For $C$ to describe a conformal motion it is necessary for its left and right quadrance to be equal and a non-zero real polynomial: $C\tilde{C}=\tilde{C}C\in\mathbb{R}[t]\setminus\{0\}$. In the case that these conditions are met we call $C$ a motion polynomial. Indeed, it describes a conformal motion as $C(t)$ is a conformal displacement for any $t\in\mathbb{R}$ with at most finitely many exceptions, namely the real roots of the quadrance polynomial $C\tilde{C}$. Since the trajectories of all points are rational curves we speak of a rational motion.
@@ 2.3 Decomposition into Simple Motions
    For generic polynomials $C$ of degree $n$ the number of factorizations depends upon the number of real roots of $C\tilde{C}$ and ranges from $n!$ for no real roots to $\frac{(2n)!}{2^{n}}$ for $2n$ roots.
@@
    tht-h is a right factor of $C$ if and only if $h$ is a right root of $C$
@@
For this we will henceforth assume $C$ to have an invertible leading coefficient. Because of $C\tilde{C}\neq 0$ this is no loss of generality as it can be ensured by a suitable rational re-parametrization. But then we can also assume that $C$ is monic, i.e. has leading coefficient $1$, as the leading coefficient can be factored out.
Since tht-h being a right factor is equivalent to $h$ being a right root of $C$, the question of finding right factors reduces to the question of finding right roots. Since a right root of the polynomial is also necessarily a right root of the quadrance of the polynomial, we can search for roots of $C\tilde{C}$. Each right factor tht-h and therefore also each right root corresponds to a monic, quadratic factor $M\coloneqq(t-h)(t-h){\,\tilde{}}$ of the quadrance polynomial. Using the Euclidean algorithm [10], we can now divide $C$ by $M$ and get
@@
for suitable polynomials $Q$, $R$ with $\deg R<\deg M=2$. This is possible because the leading coefficient $1$ of the divisor tht-h is invertible. Since we know that $h$ is a root of $C$ and $M$ it follows that it also has to be a root of $R$ [1, Lemma 1]. Furthermore, we can write $R=r_{1}t+r_{0}$, where $r_{0},r_{1}\in\mathrm{CGA}_{+}$. The question of factorization is closely related to the roots of the linear polynomial $R$. Assuming for the moment that $r_{1}$ is invertible we get the unique solution $h=-r_{1}^{-1}r_{0}$ for our particular choice of $M$. Once we have found this right factor we can divide $C$ by it and get $C^{\prime}$ of a lower degree and can iterate. Note that there is also a “left” variant of this factorization algorithm.
If the leading coefficient $r_{1}$ is not invertible, we potentially get no or infinite roots of $R$. To solve for $h$ we convert $R(h)=0$ into a system of linear equations via the coefficient-vectors with regards to the basis of $\mathrm{CGA}$. The resulting system of linear equations has no or an infinite amount of solutions which may lead to no, or infinitely many but also to finitely many factors of $C$ that can be determined by imposing further necessary condition (cf. Section 3.3). In order to capture this special situation, we define
@@
We call a factorization $C=(t-h_{1})(t-h_{2})\cdots(t-h_{n})$ of a motion polynomial $C$ irregular, if there exists an index $\ell\in\{1,2,\ldots,n\}$ such that the linear remainder polynomial $R$ obtained by dividing either
@@
[…]
@@
Definition 2.1 covers both the left and the right version of the factorization algorithm. It is, however, not clear whether an irregular “right” factorization implies an irregular “left” factorization or not. For quadratic motion polynomials however, as the remainder polynomials have the same leading coefficient, both notions coincide.